OpenStudy (anonymous):

A water balloon is tossed into the air with an upward velocity of 25 ft/s. Its height h(t) in ft after t seconds is given by the function h(t) = − 16t2 + 25t + 3. Show your work. 19. After how many seconds will the balloon hit the ground? (hint: Use the quadratic formula)

OpenStudy (anonymous):

Do you know how to use the quadratic formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

help please

OpenStudy (calculusfunctions):

Quadratic Formula: If\[a x^{2}+bx +c =0\]then\[x =\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (calculusfunctions):

When the balloon hits the ground, h(t) = 0. Thus set the equation to zero and solver for t using the quadratic formula.

OpenStudy (anonymous):

− 16t2 + 25t + 3=0

OpenStudy (anonymous):

the second part I don't know how to do it

OpenStudy (calculusfunctions):

\[-16t ^{2}+25t +3=0\]where a = -16, b = 25 and c = 3 Substitute for a, b, and c in to the formula to calculate the time t.

OpenStudy (anonymous):

wat to do next

OpenStudy (calculusfunctions):

I just explained.

OpenStudy (anonymous):

i'm not sure help me please

OpenStudy (calculusfunctions):

You don't know how to plug numbers into a formula or is this a ploy to manipulate someone to just do your work for you? Just wondering!

OpenStudy (anonymous):

can u plug in the numbers for me and I will do it